1. Given the following information estimate the monthly expected return of a portfolio of 40% AAA and 60% BBB. Assume you purchase these shares in July 2018. Enter your answer as a percent. Do not include the % sign. Round your final answer to two decimals.

2. Given the following information estimate the standard deviation of a portfolio of 40% AAA and 60% BBB monthly returns.

Stock Price Stock Price
Month AAA BBB
Jan-18 100 75
Feb-18 120 80
Mar-18 140 95
Apr-18 110 100
May-18 115 85
Jun-18 95 90
Jul-18 100 100
Answer #1

For AAA:

Average stock price = (100 + 120 + 140 + 110 + 115 + 95 + 100) / 7 = 111.42

For BBB:

Average stock price = (75 + 80 + 95 + 100 + 85 + 90 + 100) / 7 = 89.28

If investing 100 each in AAA and BBB, in July 2018, monthly expected return is calculated as

(0.4 * 111.42) + (0.6 * 89.28) = 44.57 + 53.57 = 98.14

100 - 98.14 = 1.86

Hence the monthly expected return is -1.86